Can you help me understand how to solve this word problem?
An airplane has a total of 152 seats.
The number of coach-class seats is 5 more than six
times the number of first-class seats. How many of
each type of seat are there on the plane?

Answer by stanbon(75887)

An airplane has a total of 152 seats.
The number of coach-class seats is 5 more than six
times the number of first-class seats. How many of
each type of seat are there on the plane?
-----------
Seats Equation: f + c = 152
Seats Equation: c = 6f + 5
--------
Substitute for "c" and solve for "f":
f + (6f+5) = 152
7f = 147
f = 21 (# of 1st class seats)
-----
Solve for "c":
f + c = 152
21 + c = 152
c = 131 (# of coach-class seats)
=====================================
Cheers,
Stan H.
===============